# A Compact Implicit Neural Representation for Efficient Storage of Massive 4D Functional Magnetic Resonance Imaging

### 30 Nov 2023 v1

@@ 3.6 Overview of Network’s Workflow
Based on the above analysis, we provide a summary of the workflow of the network, as shown in Fig. 2. Firstly, we apply ICA to the fMRI data and use the obtained results to initialization the learnable matrix $C_{train}$. When the training process begins, every row vector in $C_{train}$ will be the input of $K$ Conv1d blocks respectively, and the output will be concatenated into a feature map $F$. Meanwhile, the coordinate $\boldsymbol{v}$ will be embedded to serve as the input of $K$ MLPs. The output of MLPs are single weight numbers $\boldsymbol{w}^{(i)}$ which are concatenated into the weight vector $\boldsymbol{w}$. During the process of Channel Attention, $\boldsymbol{w}$ and $F$ will be integrated and serve as the input of the Feature Fusion Block. The Feature Fusion Block contains several Down Blocks and Up Blocks respectively for downsampling and upsampling, whose structures are shown in Fig. 3. Finally, a Conv1D block will be applied to the output of Feature Fusion Block and generate the output vector $\boldsymbol{x}_{\boldsymbol{v}}$.
@@ 3.7 Model Compression
Model pruning, quantization, and entropy coding are prominent techniques in model compression[20].To speed up the training process, we do not apply model pruning in our network. Upon completion of training, we proceed to quantize the network parameters and subsequently use Huffman coding as the entropy coding method.
@@ 4.1 Implementation Details
We conducted experiments on four fMRI datasets collected for different downstream tasks: Three datasets are from OpenfMRI [^1] These three datasets were obtained from the OpenfMRI database, with accession numbers ds000007, ds000101, ds00102, respectively., an open-source repository for the free and open sharing of fMRI datasets. The fourth dataset is the widely used Haxby dataset, a pioneering study of brain pattern recognition[21], which is with extended time series and suitable for the fMRI classification task.
During the training stage, we randomly selected fMRI data from the four datasets. For the first three datasets, for easier subsequent analysis, we uniformly preprocessed the fMRI data by aligning towards a standard brain template, resulting in voxel dimensions of 64$\times$64$\times$48 and a time series length of 100. The Haxby dataset comes with a pre-matched mask, and do not perform registration which is unnecessary for the classification task. Due to its longer time series, we opted to address CUDA memory limitations by slicing the data in Haxby.
In the experiments, our MLP network was configured with 5 layers, and the frequency of the coordinate embedding was set to 10. We used JPEG to compress the mean of the data, and utilized the Adamax optimizer with an initial learning rate of 8e-4. The training epoch was set to 1500. We also utilized ICA’s output network pre-training. The compression ratio of the network can be adjusted by varying parameters such as the number of ICA components, average frame compression quality, count of the MLP parameters, number of convolution channels and network layers. All these parameters can be adjusted using stored YAML files.
@@
Method ICAINR(ours) H.264 H.265 JPEG NeRF NeRV SIREN SSF DVC
@@ 4.2 Benchmark Methods and Evaluation Metrics
To demonstrate the advantages of the proposed approach, we comprehensively compare with eight state-of-the-art (SOTA) algorithms, which can be broadly categorized into three groups, including widely used traditional compression algorithms like JPEG[55], H.264[56], and H.265 (HEVC)[50], data-driven deep learning compression algorithms such as SSF[1] and DVC[30], and implicit neural representation-based compressors, including NeRV[6], NeRF[34], and SIREN[46]. For the competitors designed for 2D or 3D data, we compressed the 4D fMRI data slice by slice or concatenated the slices into 2D or 3D data.
We used the OpenCV implementation of JPEG and the FFmpeg implementation of H.264 and H.265. We set the compression ratio by calculating the corresponding bit rate. For DVC and SSF, which are data-driven methods, we changed the compression ratio by specifying the quality parameters and fine-tuned the pre-trained models provided by their authors. For SIREN and NeRF, we set the layer numbers of MLP to 7, and for NeRV, we specified the network parameters to achieve different compression ratios.
The evaluation metrics employed in our experiments can be divided into two parts. The first part involves traditional image quality evaluation metrics PSNR and SSIM, which are compared across various compression ratios. The second part pertains to downstream tasks based on fMRI data. We selected three downstream tasks, which are all classical and widely used methods in fMRI analysis. We believe that these tasks can provide a comprehensive evaluation of compression quality and fidelity, as they quantitatively measure the preservation of various informative cues within the data, including the information about stimuli-brain, intra-brain, and fMRI signal patterns, covering a significant portion of fMRI analysis. These downstream tasks encompass:
@@
ii) Brain regions Functional Connection Analysis (FCA): The brain connectome characteristics have offered valuable insights to explain the diversity of pathological conditions and behaviors across different peoples[35], and functional connection is one of the most widely used connectome characteristic, which is defined as the correlation coefficients between the voxel waveforms of different brain areas[42]. In this task, the same dataset was used as in the FLA task, with the MSDL atlas template provided in nilearn[54]. Again, we use the mean absolute difference between the correlation matrices from the ground truth and decompressed fMRI for performance evaluation.
iii) fMRI-based Classification Task (CT): Decoding or pattern recognition techniques is a significant part of fMRI analysis[21]. With the development of machine learning, there has been growing interest in decoding fMRI data by machine learning. We use a linear classifier based on SVM[9], which is most commonly used in fMRI classification[36], to do our experiment. In this task, We employed the Haxby dataset with various kinds of images serving as stimuli. We trained the SVM with the voxel waves, to distinguish stimuli images of the house and face. And we applied 10-folder cross-validation and used the cross-validation classification accuracy of the decompressed fMRI as the evaluation metric.
@@ 4.3 Experiment Results
PSNR & SSIM.
We first validate the data fidelity after fMRI compression, and compare our approach against existing compression algorithms. Here we use PSNR and SSIM as quantitative evaluation metrics, and provide the scores of different algorithms across varying compression ratios, as plotted in Fig. 4a and Fig. 4b. Notably, under similar compression ratios, our method consistently outperforms existing SOTA algorithms at both higher and lower compression ratios. When changing the compression ratio, some algorithms fluctuates, in contrast to the remarkably stable performance of our approach.
@@
Visual Results.
The visual comparison of the decompressed data by various algorithms at a fixed compression ratio ($\sim$100$\times$) is shown in Fig. 5. Note that the actual achieved compression ratio of different algorithms differs slightly, are shown in Tab. 1, because one cannot specified the final compression ratio exactly. From the results, it is evident that despite yielding relatively high PSNR and SSIM values, many algorithms failed to deliver satisfactory visual quality. For instance, JPEG and DVC suffer from block effect[26], with noticeable fragmentation between adjacent image blocks. The INR-based algorithms, like NeRF, SIREN, and NeRV, sacrificed a considerable amount of high frequency details and thus show over-smoothness. Differently, other algorithms such as H.264 and H.265 exhibited noticeable noise, and it seems that SSF almost fail to model the fMRI data.
@@
FLA Results.
To evaluate the behaviors of our compressor and other competitors on FLA, We computed the mean and standard deviation of residue between the compressed fMRI data and ground truth FLA results, as plotted in Fig. 4c and Fig. 4d. The curves indicate that our method has the smallest residue and performance fluctuation under all compression ratios among these algorithms. Among other algorithms, NeRF can achieve results slightly inferior to ours by around 0.1 and 0.15 in terms of mean and standard deviation on average, while the performance of the remaining competitors are quite limited. Additionally, we display the visualized FLA obtained by various algorithms at a compression ratio about $\sim$100$\times$, as illustrated in Fig. 6. We can observe that our method exhibits high similarity to the ground truth, while others show significant differences. These results indicate that our method effectively preserves the activities caused by stimuli, which are one of the most important components of fMRI signals.
FCA Results.
Regarding to the influence of compression on FCA, similar to FLA, we illustrated the mean and standard deviation of residual after decompression in Fig. 4e and Fig. 4f. The curves show that our method has much less information loss than other algorithms if the compression ratio is lower than 170$\times$, and the average loss approximates to zero when compression ratio is around 30$\sim$40. When compression ratio is higher, the performance in mean residual of NeRF, SIREN, NeRV and our method is similar. From the the standard deviation, one can conclude that our approach is the lowest performance fluctuation among all the algorithms working at a similar compression ratio, and achieves stable performance (standard deviation close to 0) when the compression ratio is lower than 100$\times$. We also compare the visualized brain connectivity at a compression ratio of about 100$\times$ in Fig. 7, from which one can observe that our compressor presents the highest similarity to ground truth. These results show our capability of keeping the correlation information among brain regions.
CT Results.
To test the effectiveness of our compressor and advantageous over previous competitors in the successive classification task (CT), we calculated the 10-folder cross validation accuracy and AUC of the decompressed data compared to the original data before compression, as shown in Fig. 4g and Fig. 4h. The results demonstrate that our algorithm outperforms other methods at compression ratio lower than 170, and can achieve accuracy and AUC close to the ground truth, 93.89% and 0.9975 respectively, when the compression ratio reduces below 100$\times$. In other words, our compression preserves most of the distinctive features in the fMRI data volumes.
@@ 4.4 Ablation Studies
In this section, we quantitatively identify the contribution of the key modules in the proposed approach.
Decomposition of Neuronal Activation Patterns.
We compared SIREN, which is INR-based method modeling the data directly, with our method. We removed the Feature Fusion Block in our network to eliminate its impact. As listed in 2nd, 5th and 6th rows of Tab. 2), which showed better performance in terms of all metrics. The results show that using pattern decomposition is advantageous over directly modeling the 4D fMRI data with INR, because of the high complexity of fMRI signals and various underlying correlations.
Feature Fusion Block.
At a similar compression ratio, we compare the quality of the decompressed data before and after removing the feature fusion block. The results are shown in the 1st and 2nd rows of Tab. 2, demonstrating that the Feature Fusion Block achieves higher data fidelity of the target fMRI signals than naive linear modeling.
ICA Initialization.
To verify the effectiveness of ICA initialization, we compared the performance with the $C_{\text{train}}$ matrix uses ICA initialization, Uniform Initialization, and Normal Initialization 1st, 3rd and 4th rows in Tab. 2). The results, especially the performance in downstream tasks, show that ICA initialization can guide the network to model the signals better, especially the neural activation information and correlations.
@@ 5 Summary and Discussions
In this paper, we introduce a novel INR-based compression paradigm for fMRI data. Leveraging the strong representation capability of deep neural networks and fMRI’s unique characteristics, we compactly describe the neural activation patterns and their spatial distributions, which reduces both temporal and spatial redundancies in the raw recordings.
Advantages.
From the comprehensive experiments, our algorithm outperforms existing SOTAs in terms of data fidelity such as PSNR and SSIM, as well as evaluation metrics for downstream fMRI tasks, which meets the fidelity requirements of medical image compression. From the comprehensive experiments, our algorithm achieves a higher compression ratio and outperforms existing SOTA methods in terms of both image fidelity and downstream tasks. Notably, our study pioneers the application of INR-based compression methods to four-dimensional biomedical data. Furthermore, we present a novel INR-based compression framework, offering valuable insights for future research in INR-based biomedical image compression.
Limitations and Future Work.
As a preliminary research on INR-based fMRI compression, our approach can be extended in several aspects before becoming a mature compression tool. Firstly, as a deep learning-based compression algorithm, our approach takes longer time than conventional compressors. Therefore, we are exploring novel architectures and techniques, such as meta-learning, for fast model learning. Secondly, we use all the initial components obtained from ICA, but some of them contain little information about the neural activities. We are also working on strategies for extracting informative signals to avoid compressing unnecessary components and thus build a more compact compression. Thirdly, we adopt a widely used method for network compression, and there is room for designing more tailored model compression techniques for our network in the future.

### 29 Feb 2024 v2

@@ 3.8 Overview of Network’s Workflow
Based on the above analysis, we provide a summary of the workflow of the network, as shown in Fig. 2. We apply ICA and train an individual network for each fMRI data to get higher fidelity. Firstly, we calculate the mean frame of the fMRI data and subtract it from each frame. We will store the mean frame individually utilizing conventional compression like PNG due to the high structure similarities between frames. Before training, we apply ICA to the fMRI data in which the black areas are masked. The ICA results will be used to initialize $C_{train}$ and pre-train the network. When the training process begins, every row vector in $C_{train}$ will be the input of $K$ Conv1d blocks respectively, and the output will be concatenated into a feature map $F$. Meanwhile, the coordinate $\boldsymbol{v}$ will be embedded to serve as the input of $K$ MLPs, and the embedding method we leverage is the same one in NeRF[34]. The output of MLPs are single weight numbers $\boldsymbol{w}^{(i)}$ which are concatenated into the weight vector $\boldsymbol{w}$. During the process of Channel Attention, $\boldsymbol{w}$ and $F$ will be integrated and serve as the input of the Feature Fusion Block. The Feature Fusion Block contains several Down Blocks and Up Blocks respectively for downsampling and upsampling, whose structures are shown in Fig. 4. Finally, a Conv1D block will be applied to the output of Feature Fusion Block and generate the output vector $\boldsymbol{x}_{\boldsymbol{v}}$.
@@ 3.9 Model Compression
Model pruning, quantization, and entropy coding are prominent techniques in model compression[20].To speed up the training process, we do not apply model pruning in our network. Upon completion of training, we proceed to quantize the network parameters and subsequently utilize Huffman coding as the entropy coding method.
@@ 4.1 Implementation Details
We conducted experiments on four fMRI datasets collected for different downstream tasks: Three datasets are from OpenfMRI [^1] These three datasets were obtained from the OpenfMRI database, with accession numbers ds000007, ds000101, ds00102, respectively., an open-source repository for the free and open sharing of fMRI datasets. The fourth dataset is the widely used Haxby dataset, a pioneering study of brain pattern recognition[21], which has extended time series and suitable for the fMRI classification task.
During the training stage, we randomly selected fMRI data from the four datasets. For the first three datasets, in order to facilitate subsequent analysis, we uniformly preprocessed the fMRI data by aligning it to a standard brain template. This resulted in voxel dimensions of 64$\times$64$\times$48 and a time series length of 100. The Haxby dataset comes with a pre-matched mask. And we did not perform registration, as it is unnecessary for the classification task. Due to its longer time series, we addressed CUDA memory limitations by slicing the data in Haxby.
In the experiments, our MLP network was configured with five layers, and the frequency of the coordinate embedding was set to 10. We used JPEG to compress the mean of the data, and utilized the Adamax optimizer with an initial learning rate of 8e-4. The training epoch was set to 1500. We also utilized ICA’s output network pre-training. The compression ratio of the network can be adjusted by varying parameters such as the number of ICA components, average frame compression quality, count of the MLP parameters, number of convolution channels and network layers. All these parameters can be adjusted using stored YAML files.
@@
Method ICAINR(ours) H.264 H.265 JPEG NeRF HNeRV SIREN SSF DVC
@@ 4.2 Benchmark Methods and Evaluation Metrics
To demonstrate the advantages of the proposed approach, we comprehensively compare it with eight state-of-the-art (SOTA) algorithms, which can be broadly categorized into three groups: widely used traditional compression algorithms like JPEG[55], H.264[56], and H.265 (HEVC)[50], data-driven deep learning compression algorithms such as SSF[1] and DVC[30], and implicit neural representation-based compressors, including HNeRV[5], NeRF[34], and SIREN[46]. For the competitors designed for 2D or 3D data, we compressed the 4D fMRI data slice by slice or concatenated the slices into 2D or 3D data.
We utilized the OpenCV implementation of JPEG and the FFmpeg implementation of H.264 and H.265. We set the compression ratio by calculating the corresponding bit rate. For DVC and SSF, which are data-driven methods, we changed the compression ratio by specifying the quality parameters and fine-tuning the pre-trained models provided by their authors. For SIREN and NeRF, we set the layer numbers of MLP to 7, and for HNeRV, we specified the network parameters to achieve different compression ratios.
The evaluation metrics employed in our experiments can be divided into two parts. The first part involves traditional image quality evaluation metrics PSNR and SSIM, which are compared across various compression ratios. The second part pertains to downstream tasks based on fMRI data. We selected three downstream tasks, which are all classical and widely used methods in fMRI analysis to comprehensively evaluate the compression quality and fidelity. These tasks quantitatively measure the preservation of various informative cues within the data, including the information about stimuli-brain, intra-brain, and fMRI signal patterns, covering a significant portion of fMRI analysis. These downstream tasks encompass the following:
@@
ii) Brain regions Functional Connectivity Analysis (FCA): The brain connectome characteristics have offered valuable insights to explain the diversity of pathological conditions and behaviors across different peoples[35], and functional connectivity is one of the most widely used connectome characteristics. It is defined as the correlation coefficients between the voxel waveforms of different brain areas[42]. In this task, the same dataset was used as in the FLA task, with the MSDL atlas template provided in nilearn[54]. Again, we use the mean absolute difference between the correlation matrices from the ground truth and decompressed fMRI for performance evaluation.
iii) fMRI-based Classification Task (CT): Decoding or pattern recognition techniques are a significant part of fMRI analysis[21]. With the development of machine learning, there has been growing interest in decoding fMRI data by machine learning. We use a linear classifier based on SVM[9], which is most commonly used in fMRI classification[36], to do our experiment. In this task, we employed the Haxby dataset with various kinds of images serving as stimuli. We trained the SVM with the voxel waves, to distinguish stimuli images of the house and face. We applied 10-fold cross-validation and used the cross-validation classification accuracy of the decompressed fMRI as the evaluation metric.
@@ 4.3 Experiment Results
PSNR & SSIM.We first validate the data fidelity after fMRI compression, and compare our approach against existing compression algorithms. Here we use PSNR and SSIM as quantitative evaluation metrics, and provide the scores of different algorithms across varying compression ratios, as plotted in Fig. 5a and Fig. 5b. Notably, under similar compression ratios, our method consistently outperforms existing SOTA algorithms at both higher and lower compression ratios. When changing the compression ratio, some algorithms fluctuate, in contrast to the remarkably stable performance of our approach.
@@
Visual Results.The visual comparison of the decompressed data by various algorithms at a fixed compression ratio ($\sim$100$\times$) is shown in Fig. 6. Note that the actual achieved compression ratio of different algorithms differs slightly, as shown in Tab. 1, because one cannot specify the final compression ratio exactly. From the results, it is evident that despite yielding relatively high PSNR and SSIM values, many algorithms failed to deliver satisfactory visual quality. For instance, JPEG and DVC suffer from block effect[26], with noticeable fragmentation between adjacent image blocks. The INR-based algorithms, like NeRF, SIREN, and HNeRV, sacrificed a considerable amount of high-frequency details and thus show over-smoothness. In contrast, other algorithms such as H.264 and H.265 exhibited noticeable noise, and it seems that SSF almost fails to model the fMRI data.
@@
FLA Results.To evaluate the behaviors of our compressor and other competitors on FLA, we computed the mean and standard deviation of residue between the compressed fMRI data and ground truth FLA results, as plotted in Fig. 5c and Fig. 5d. The curves indicate that our method has the smallest residue and performance fluctuation under all compression ratios among these algorithms. Among other algorithms, NeRF can achieve results slightly inferior to ours by around 0.1 and 0.15 in terms of mean and standard deviation on average, while the performance of the remaining competitors is quite limited. Additionally, we display the visualized FLA obtained by various algorithms at a compression ratio about $\sim$100$\times$, as illustrated in Fig. 7. We can observe that our method exhibits high similarity to the ground truth, while others show significant differences. These results indicate that our method effectively preserves the activities caused by stimuli, which are one of the most important components of fMRI signals.
FCA Results.Regarding to the influence of compression on FCA, similar to FLA, we illustrated the mean and standard deviation of residual after decompression in Fig. 5e and Fig. 5f. The curves show that our method has much less information loss than other algorithms if the compression ratio is lower than 170$\times$, and the average loss approximates zero when compression ratio is around 30$\sim$40. When compression ratio is higher, the performance in mean residual of NeRF, SIREN, and our method is similar. From the the standard deviation, one can conclude that our approach exhibits the lowest performance fluctuation among all the algorithms working at a similar compression ratio, and achieves stable performance (standard deviation close to 0) when the compression ratio is lower than 100$\times$. We also compare the visualized brain connectivity at a compression ratio of about 100$\times$ in Fig. 8, from which one can observe that our compressor presents the highest similarity to ground truth. These results demonstrate our capability to preserve the correlation information among brain regions.
CT Results.To test the effectiveness of our compressor and its advantage over previous competitors in the successive classification task (CT), we calculated the 10-fold cross validation accuracy and AUC of the decompressed data compared to the original data before compression, as shown in Fig. 5g and Fig. 5h. The results demonstrate that our algorithm outperforms other methods at compression ratio lower than 170, and can achieve accuracy and AUC close to the ground truth, 93.89% and 0.9975 respectively, when the compression ratio falls below 100$\times$. In other words, our compression preserves most of the distinctive features in the fMRI data volumes.
@@ 4.4 Ablation Studies
In this section, we quantitatively determine the contribution of the key modules in the proposed approach.
Decomposition of Neuronal Activation Patterns. We compared SIREN, an INR-based method modeling the data directly, with our method. We removed the Feature Fusion Block in our network to eliminate its impact. As listed in 2nd, 5th and 6th rows of Tab. 2), which showed better performance in terms of all metrics. The results show that using pattern decomposition is advantageous over directly modeling the 4D fMRI data with INR, given the high complexity of fMRI signals and various underlying correlations.
Feature Fusion Block. At a similar compression ratio, we compare the quality of the decompressed data before and after removing the feature fusion block. The results are shown in the 1st and 2nd rows of Tab. 2, demonstrating that the Feature Fusion Block achieves higher data fidelity of the target fMRI signals than naive linear modeling.
ICA Initialization. To verify the effectiveness of ICA initialization, we compared the performance with the $C_{\text{train}}$ matrix uses ICA initialization, Uniform Initialization, and Normal Initialization 1st, 3rd and 4th rows in Tab. 2). The results, especially the performance in downstream tasks, demonstrate that ICA initialization can enhance the network’s ability to model the signals, especially the neural activation information and correlations.
@@ 5 Summary and Discussions
In this paper, we introduce a novel INR-based compression paradigm for fMRI data. By leveraging the strong representation capability of deep neural networks and taking into account of fMRI’s unique characteristics, we aim to compactly describe the neural activation patterns and their spatial distributions, which reduces both temporal and spatial redundancies in the raw recordings.
Advantages. Based on the comprehensive experiments conducted, our algorithm outperforms existing SOTAs in terms of data fidelity such as PSNR and SSIM, as well as evaluation metrics for downstream fMRI tasks, which meets the fidelity requirements for medical image compression. Moreover, our algorithm achieves a higher compression ratio and outperforms existing SOTA methods in terms of both image fidelity and downstream tasks. Notably, our study pioneers the application of INR-based compression methods to four-dimensional biomedical data. Additionally, we present a novel INR-based compression framework, offering valuable insights for future research in INR-based biomedical image compression.
Limitations and Future Work. As a preliminary research on INR-based fMRI compression, our approach can be extended in several aspects before becoming a mature compression tool. Firstly, as a deep learning-based compression algorithm, our approach takes longer time than conventional compressors. Therefore, we are exploring novel architectures and techniques, such as meta-learning, for fast model learning. Secondly, we use all the initial components obtained from ICA, but some of them contain little information about the neural activities. We are also working on strategies for extracting informative signals to avoid compressing unnecessary components and thus build a more compact compression. Thirdly, we adopt a widely used method for network compression, and there is room for designing more tailored model compression techniques for our network in the future.